OpenStudy (rivershinesdragon):

Help: y=x^2 x=z+y z=4x+y Solve x+y+z =?

OpenStudy (faiqraees):

In terms of what should we give our answer?

OpenStudy (rivershinesdragon):

May you please explain what you mean by terms? If you mean something like form, I think it's supposed to be simplified as far as possible, whether there are variables left over or no.

OpenStudy (rivershinesdragon):

I guessed 2x^2+8x+2y, but I'm not sure...

OpenStudy (faiqraees):

In terms of what means, what variables should our expression contain Example 4xy (given in terms of x and y) 4x^2 (given in terms of x) 4yz (given in terms of y and z)

OpenStudy (phi):

terms are formed by numbers/letters multiplied together (no add or subtract) adding y=x^2 x=z+y z=4x+y is , in some sense, simple: you put a "+" sign between each term to show them being added together: x+y+z = x^2 + y+z + 4x+y However, they usually mean add, and simplify as much as possible. in this case we have 2 y's y+y which is 2y x+y+z = x^2 + 4x + 2y + z

OpenStudy (mathmate):

This is a system of homogeneous equations, where there is the trivial answer of x=0,y=0,z=0. However, there is also a set of non-trivial solutions. If you can solve for the numerical values of x,y and z, then you can easily find x+y+z. You can solve by substitution or a combination of any other techniques.

OpenStudy (mathmate):

Hint: the non-trivial solution has all rational values, y>0 (obviously), and the other two are negative fractions. The sum x+y+z is a negative integer.